# In search of maximal branes on hyper-Kähler manifolds

### v1

@@ 7.1. Holomorphic symplectic involutions
Here we treat (BBB)-branes and deduce Theorem˜1.4 from results obtained in Section˜6.
Proof of Theorem˜1.4.
Thanks to [30, Theorem 3.2], for any symplectic involution $\tau$ on a hyper-Kähler manifold $X$ of K3[n] type, the pair $(X,\tau)$ is deformation equivalent to a pair $(S^{[n]},\sigma^{[n]})$ for a K3 surface $S$ and natural involution $\sigma^{[n]}$ induced by a symplectic involution $\sigma$ on $S$. As a consequence, the fixed locus of $\tau$ is diffeomorphic to the fixed locus of $\sigma^{[n]}$, which is never maximal by ˜6.6, unless the involution is trivial. ∎
@@ 7.2. Anti-symplectic involutions
The goal of this section is to prove Theorem˜1.2.
@@
Most of the statements can be deduced from [6, Proposition 2.8]. As our case is simplified thanks to the assumption (7.2) (which is amount to the condition $a=0$ in loc. cit.), we give a direct proof for convenience of the reader.
@@
We denote by $\overline{\sigma}$ the action of $\sigma$ on the discriminant group $A_{L}\cong\mathbb{Z}/(2n-2)\mathbb{Z}$. By Markman [41, Lemma 9.2],
@@
    If $L^{\sigma}$ is unimodular, since it is even and has index (1,)(1,-), by classification of such lattices (see [43, Chapter II]), it is determined by its rank and signature, and the signature is divisible by 8. Hence
@@
    with $i=0,1$ or 2. Consequently, $L^{\sigma-}$ is even, of index (2,)(2,-), with length $\ell(A_{L^{\sigma-}})=\ell(\mathbb{Z}/(2n-2)\mathbb{Z})=1$ and rank $5,13$ or 21. By Nikulin [47, Corollary 1.13.3], $L^{\sigma-}$ is determined by its rank, index and discriminant form. Then the classification follows.
@@
    with $i=0,1$ or 2. Hence $L^{\sigma}$ is even, indefinite, with length 1, and with rank at least 3. Applying Nikulin [47, Corollary 1.13.3], we conclude the classification.
Before proceeding further with the proof, we recall the following generalization of Eichler’s criterion (see [13, §10] for the classical version). For a vector $v$ in a lattice , we denote by $v^{*}$ the class $\frac{v}{\mathrm{div}(v)}$ in the discriminant group $A$, where $\mathrm{div}(v)$ is the divisibility of $v$.
Theorem 7.2 ([21, Proposition 3.3]).
@@
Let us denote by $L^{\prime}$ the sublattice $L^{\sigma}$ or $L^{\sigma-}$, depending on $L^{\sigma-}$ or $L^{\sigma}$ is unimodular. From the classification in Lemma˜7.1, it is clear that there exists a primitive element $x\in L^{\prime}$ with $x^{2}=2-2n$ and $\mathrm{div}(x)=2n-2$ such that $A_{L}$ is generated by $x^{*}:=\frac{x}{\mathrm{div}(x)}$. On the other hand, by hypothesis $A_{L}$ is also generated by $v^{*}$, hence there exists an integer $k$ that is coprime to 2n22n-2, such that $v^{*}=kx^{*}$ in $A_{L}$. Therefore, we have the following equality in $\mathbb{Q}/2\mathbb{Z}$, since the lattice $L$ is even:
@@
This readily implies that $\frac{k^{2}-1}{2n-2}\in 2\mathbb{Z}$. Again by the classification in Lemma˜7.1, there is always a copy of the hyperbolic plane $U$ in $x^{\perp_{L^{\prime}}}$ as a direct summand. As the hyperbolic plane quadratic form represents all even integers, there exists $w\in L^{\prime}$ such that $w\perp x$ and $w^{2}=\frac{k^{2}-1}{2n-2}$.
@@
For any K3 surface $S$, we have that $X$ is deformation equivalent to $S^{[n]}$ and hence there is a parallel transport operator inducing an isometry $f\colon H^{2}(X,\mathbb{Z})\cong H^{2}(S^{[n]},\mathbb{Z})$. Consider the class $\delta\in H^{2}(S^{[n]},\mathbb{Z})$ corresponding to $\frac{1}{2}[E]$ with $E$ the exceptional divisor of $S^{[n]}$. The element $f^{-1}(\delta)^{*}$ generates the discriminant group of $L=H^{2}(X,\mathbb{Z})$, hence by Lemma˜7.3 there exists an element $\epsilon\in L^{\sigma-}$, or $\epsilon\in L^{\sigma}$ respectively, with the same square and divisibility as $f^{-1}(\delta)$ and such that $\epsilon^{*}=f^{-1}(\delta)^{*}$ in $A_{H^{2}(X,\mathbb{Z})}$.
Therefore, by Theorem˜7.2, we can find an orientation-preserving isometry acting trivially on the discriminant group $g\in\widetilde{O}^{+}(L)$ such that $g(\epsilon)=f^{-1}(\delta)$. Markman proved that $\operatorname{Mon}^{2}(X)\cong W(H^{2}(X,\mathbb{Z})):=\{h\in O^{+}(H^{2}(X,\mathbb{Z}))~|~\bar{h}=\pm\operatorname{id}\in O(A_{H^{2}(X,\mathbb{Z})})\}$ if $X$ is of K3[n]-type, see for example [41]. This means that the isometry $g$ is induced by parallel transport and hence it is also the case of the composition $\phi:=f\circ g\colon H^{2}(X,\mathbb{Z})\to H^{2}(S^{[n]},\mathbb{Z})$, which is then the restriction of a ring isomorphism that we call again $\phi\colon H^{*}(X,\mathbb{Z})\xrightarrow{\cong}H^{*}(S^{[n]},\mathbb{Z})$. ∎
Proof of Theorem˜1.2.
Keep the notation as before. We separate the proof according to the four cases in Lemma˜7.1, firstly in Case 1 or 2, then in Case 3 or 4.
In Case 1 or Case 2 of Lemma˜7.1: then $L^{\sigma}$ or $L^{\sigma-}$ is unimodular. By Lemma˜7.3 and Lemma˜7.4, we can find a graded ring isomorphism $\phi\colon H^{*}(X,\mathbb{Z})\cong H^{*}(S^{[n]},\mathbb{Z})$ sending $\epsilon$ to $\delta$. We set
@@
From the classification in Case 1 and Case 2 in Lemma˜7.1, we observe that $\epsilon^{\perp}\cap H^{2}(X,\mathbb{Z})^{\sigma}\neq 0$ or respectively $\epsilon^{\perp}\cap H^{2}(X,\mathbb{Z})^{\sigma-}\neq 0$, so that we are able to pick a non-zero primitive element $\alpha\in\delta^{\perp}\cap H^{2}(S^{[n]},\mathbb{Z})^{\iota-}$ when $\iota(\delta)=\delta$ (Case 2), or $\alpha\in\delta^{\perp}\cap H^{2}(S^{[n]},\mathbb{Z})^{\iota}$ when $\iota(\delta)=-\delta$ (Case 1). Extend $\alpha$ to a basis of $H^{2}(S,\mathbb{Z})$. Consider the basis element
@@
from Theorem 3.1 and the element
@@
In Case 3 of Lemma˜7.1 (this can only happen when $n\geq 4$ is an even integer): then $A_{L^{\sigma}}\cong\mathbb{Z}/2\mathbb{Z}$ and $A_{L^{\sigma-}}\cong\mathbb{Z}/{(n-1)\mathbb{Z}}$, hence $\tau(\sigma)=-1$. We will use notations as in Section˜4.4. In particular,
@@
Let us choose one $U$-summand in $Q(X,\mathbb{Z})\cong E_{8}(-1)^{\oplus 2}\oplus U^{\oplus 4}$ and pick in it two orthogonal primitive elements $u,v$ such that $u^{2}=2n-2$, $v^{2}=2-2n$ and $u+v$ is divisible by 2n22n-2. According to [58, Theorem 6], there exists an isometry $\phi:Q(X,\mathbb{Z})\to Q(X,\mathbb{Z})$ that sends $u$ to $\gamma$. We put $\delta^{\prime}=\phi(v)$ and define $w:=\frac{\delta^{\prime}+\gamma}{2n-2}\in Q(X,\mathbb{Z})$. In particular, $\delta^{\prime 2}=2-2n$, $\delta^{\prime}\in L^{\prime}$ and the divisibility of $\delta^{\prime}$ in $L^{\prime}$ is 2n22n-2.
@@
For Case 4 of Lemma˜7.1, the proof is the same as in Case 3. The only difference is that now $A_{L^{\sigma}}\cong\mathbb{Z}/(n-1)\mathbb{Z}$ and $A_{L^{\sigma-}}\cong\mathbb{Z}/{2\mathbb{Z}}$, so we have $\tau(\sigma)=1$. Again by Theorem˜4.2, we have the following (instead of (7.8)):
@@ 7.3. Anti-holomorphic involutions
Proof of ˜1.3.
Since maximality is a topological property, we can use hyper-Kähler rotation to change the complex structure. More precisely, let $X$ be a hyper-Kähler manifold of K3[n]-type ($n\geq 2$), and let $\sigma$ be a real structure with respect to the original complex structure $I$. By ˜4.1, there is another complex structure $K$ on $X$ such that $\sigma$ is holomorphic and anti-symplectic. Applying Theorem˜1.2 to the anti-symplectic holomorphic involution $\sigma$ on $(X,K)$, we deduce that $\sigma$ is not maximal. ∎
@@ 8.1. Generalization beyond biregular involutions
By inspecting the proof of Theorem˜1.2, one sees that what is actually proved is the following more general result.
@@
The only place in the proof of Theorem˜1.2 where we used that $\sigma$ is a geometric involution is at the beginning of the proof of Lemma˜7.1, where we argued that the signature of $H^{2}(X,\mathbb{Z})^{\sigma}$ is $(\geq 1,-)$ since $\sigma$ preserves a Kähler class. But the claim for the signature holds in general for any Hodge monodromy operator: since $\sigma\in\operatorname{Mon}^{2}(X)$ is of trivial spinor norm, it preserves $\mathcal{C}^{\circ}_{X}$, the connected component of the positive cone containing Kähler classes. Since $\mathcal{C}^{\circ}_{X}$ is convex, for any Kähler class $\omega$, the class $\frac{\omega+\sigma^{*}(\omega)}{2}$ is $\sigma$-invariant with positive square. Hence the signature of $H^{2}(X,\mathbb{Z})^{\sigma}$ is $(\geq 1,-)$. The rest of the proof of Lemma˜7.1 goes through verbatim. ∎
In addition to Theorem˜1.2 as a special case, Theorem˜8.1 can be applied more generally to birational automorphisms, since a birational automorphism on a compact hyper-Kähler manifold induces a Hodge monodromy operator on its cohomology. Consequently, Theorem˜8.1 implies the following result:
@@ 8.2. Hyper-Kähler varieties of other deformation types
In general, do there exist maximal branes on compact hyper-Kähler manifolds of dimension >2>2. How about the other known deformation types e.g. Kumn-type, OG10-type and OG6-type?
@@
8.3. (BBB)-branes on compact hyper-Kähler manifolds
Theorem˜1.2 and ˜1.3 can be reinterpreted as the claim that the Smith inequality Theorem˜1.1 is not optimal for anti-symplectic or anti-holomorphic involutions on compact hyper-Kähler manifolds of K3[n]-type for $n\geq 2$.
Of course, Theorem˜1.4 can be reformulated similarly, but we want to argue that the upper bound given by the Smith inequality for holomorphic symplectic involutions on compact hyper-Kähler manifolds is presumably very far from optimal. Already for K3 surfaces, any symplectic involution has 8 isolated fixed points, whose total $\operatorname{\mathbb{F}_{2}}2$-Betti number is 8, which is much smaller than the one imposed by the Smith inequality (24 in this case).
In fact, more generally, the fixed loci of symplectic involutions on compact hyper-Kähler manifolds of all known deformation types have been computed. Let $X$ be a compact hyper-Kähler manifold equipped with a holomorphic symplectic involution $\sigma$.
  • If $X$ is of K3[n]-type, by [30, Theorem 1.1], the fixed loci of $\sigma$ is the disjoint union of $\sumop\displaylimits_{2m=n-k-2l}{8\choose k}{k\choose l}$ copies of hyper-Kähler manifolds of K3[m]-type with $m$ running through integers from $\max(0,\lceil\frac{n}{2}\rceil-12)$ to $\lfloor\frac{n}{2}\rfloor$. A computation in the first few values of $n$ shows that the total Betti number of the fixed locus tends to be much smaller than the total Betti number of $X$. For example, when $n=2$ (resp. $n=3$), the fixed locus consists of a K3 surface and 28 points (resp. 8 K3 surfaces and 64 points), whose total Betti number is 52 (resp. 256), while $X$ has total Betti number 324 (resp. 3200).
  • If $X$ is of $\operatorname{Kum}_{n}$-type, by [30, Theorem 1.3], the fixed loci of $\sigma$ is the disjoint union of $N_{n+1,m}$ copies of hyper-Kähler manifolds of K3[m]-type with $m$ running through integers from $\max(0,\lceil\frac{n+1}{2}\rceil-24)$ to $\lfloor\frac{n+1}{2}\rfloor$, where the number $N_{n,m}$ is defined before [30, Theorem 3.9]. Similar to the previous case, a computation in the first few values of $n$ shows that the Smith inequality tends to be very far from optimal.
  • If $X$ is of OG6-type, by [22, Theorem 1.1], $\sigma$ acts trivially on the second cohomology. Such involutions are classified in [44, Theorem 5.2], and their fixed loci are computed in [44, Propositions 6.1, 6.3, 6.7]: it is either 16 K3 surfaces, or 16 points, or 2 K3 surfaces. Hence the maximal total $\operatorname{\mathbb{F}_{2}}2$-Betti number of the fixed locus is at most 384, while the total $\operatorname{\mathbb{F}_{2}}2$-Betti number of $X$ is 1920.
  • If $X$ is of OG10-type, by [19, Theorem 1.1], there is no non-trivial finite-order symplectic automorphism.
Question 8.6.

### v2

@@ 7.1. Holomorphic symplectic involutions
Here we treat (BBB)-brane involutions and deduce Theorem˜1.6 from results obtained in Section˜6.
Proof of Theorem˜1.6.
Thanks to [35, Theorem 1.0.3 and Theorem 2.0.6] [^7] One can also directly use [34, Corollary 3.3], whose proof relies on [34, Theorem 3.2]. Although the statement of [34, Theorem 3.2] for Kummer type hyper-Kähler varieties does not hold, it is correct in the case of K3[n]-type., for any symplectic involution $\tau$ on a hyper-Kähler manifold $X$ of K3[n] type, the pair $(X,\tau)$ is deformation equivalent to a pair $(S^{[n]},\sigma^{[n]})$ for a K3 surface $S$ and natural involution $\sigma^{[n]}$ induced by a symplectic involution $\sigma$ on $S$. As a consequence, the fixed locus of $\tau$ is diffeomorphic to the fixed locus of $\sigma^{[n]}$, which is never maximal by ˜6.6, unless the involution is trivial. ∎
Remark 7.1.
In Section˜8.3, we give an alternative proof of this result which also covers the case of hyper-Kähler manifolds of OG6 type; see ˜8.9.
@@ 7.2. Anti-symplectic involutions
The goal of this section is to prove Theorem˜1.4.
@@
Most of the statements can be deduced from [7, Proposition 2.8]. As our case is simplified thanks to the assumption (7.2) (which is amount to the condition $a=0$ in loc. cit.), we give a direct proof for convenience of the reader.
@@
We denote by $\overline{\sigma}$ the action of $\sigma$ on the discriminant group $A_{L}\cong\mathbb{Z}/(2n-2)\mathbb{Z}$. By Markman [47, Lemma 9.2],
@@
    If $L^{\sigma}$ is unimodular, since it is even and has index (1,)(1,-), by classification of such lattices (see [49, Chapter II]), it is determined by its rank and signature, and the signature is divisible by 8. Hence
@@
    with $i=0,1$ or 2. Consequently, $L^{\sigma-}$ is even, of index (2,)(2,-), with length $\ell(A_{L^{\sigma-}})=\ell(\mathbb{Z}/(2n-2)\mathbb{Z})=1$ and rank $5,13$ or 21. By Nikulin [54, Corollary 1.13.3], $L^{\sigma-}$ is determined by its rank, index and discriminant form. Then the classification follows.
@@
    with $i=0,1$ or 2. Hence $L^{\sigma}$ is even, indefinite, with length 1, and with rank at least 3. Applying Nikulin [54, Corollary 1.13.3], we conclude the classification.
Before proceeding further with the proof, we recall the following generalization of Eichler’s criterion (see [15, §10] for the classical version). For a vector $v$ in a lattice , we denote by $v^{*}$ the class $\frac{v}{\mathrm{div}(v)}$ in the discriminant group $A$, where $\mathrm{div}(v)$ is the divisibility of $v$.
Theorem 7.3 ([24, Proposition 3.3]).
@@
Let us denote by $L^{\prime}$ the sublattice $L^{\sigma}$ or $L^{\sigma-}$, depending on $L^{\sigma-}$ or $L^{\sigma}$ is unimodular. From the classification in Lemma˜7.2, it is clear that there exists a primitive element $x\in L^{\prime}$ with $x^{2}=2-2n$ and $\mathrm{div}(x)=2n-2$ such that $A_{L}$ is generated by $x^{*}:=\frac{x}{\mathrm{div}(x)}$. On the other hand, by hypothesis $A_{L}$ is also generated by $v^{*}$, hence there exists an integer $k$ that is coprime to 2n22n-2, such that $v^{*}=kx^{*}$ in $A_{L}$. Therefore, we have the following equality in $\mathbb{Q}/2\mathbb{Z}$, since the lattice $L$ is even:
@@
This readily implies that $\frac{k^{2}-1}{2n-2}\in 2\mathbb{Z}$. Again by the classification in Lemma˜7.2, there is always a copy of the hyperbolic plane $U$ in $x^{\perp_{L^{\prime}}}$ as a direct summand. As the hyperbolic plane quadratic form represents all even integers, there exists $w\in L^{\prime}$ such that $w\perp x$ and $w^{2}=\frac{k^{2}-1}{2n-2}$.
@@
For any K3 surface $S$, we have that $X$ is deformation equivalent to $S^{[n]}$ and hence there is a parallel transport operator inducing an isometry $f\colon H^{2}(X,\mathbb{Z})\cong H^{2}(S^{[n]},\mathbb{Z})$. Consider the class $\delta\in H^{2}(S^{[n]},\mathbb{Z})$ corresponding to $\frac{1}{2}[E]$ with $E$ the exceptional divisor of $S^{[n]}$. The element $f^{-1}(\delta)^{*}$ generates the discriminant group of $L=H^{2}(X,\mathbb{Z})$, hence by Lemma˜7.4 there exists an element $\epsilon\in L^{\sigma-}$, or $\epsilon\in L^{\sigma}$ respectively, with the same square and divisibility as $f^{-1}(\delta)$ and such that $\epsilon^{*}=f^{-1}(\delta)^{*}$ in $A_{H^{2}(X,\mathbb{Z})}$.
Therefore, by Theorem˜7.3, we can find an orientation-preserving isometry acting trivially on the discriminant group $g\in\widetilde{O}^{+}(L)$ such that $g(\epsilon)=f^{-1}(\delta)$. Markman proved that $\operatorname{Mon}^{2}(X)\cong W(H^{2}(X,\mathbb{Z})):=\{h\in O^{+}(H^{2}(X,\mathbb{Z}))~|~\bar{h}=\pm\operatorname{id}\in O(A_{H^{2}(X,\mathbb{Z})})\}$ if $X$ is of K3[n]-type, see for example [47]. This means that the isometry $g$ is induced by parallel transport and hence it is also the case of the composition $\phi:=f\circ g\colon H^{2}(X,\mathbb{Z})\to H^{2}(S^{[n]},\mathbb{Z})$, which is then the restriction of a ring isomorphism that we call again $\phi\colon H^{*}(X,\mathbb{Z})\xrightarrow{\cong}H^{*}(S^{[n]},\mathbb{Z})$. ∎
Proof of Theorem˜1.4.
Keep the notation as before. We separate the proof according to the four cases in Lemma˜7.2, firstly in Case 1 or 2, then in Case 3 or 4.
In Case 1 or Case 2 of Lemma˜7.2: then $L^{\sigma}$ or $L^{\sigma-}$ is unimodular. By Lemma˜7.4 and Lemma˜7.5, we can find a graded ring isomorphism $\phi\colon H^{*}(X,\mathbb{Z})\cong H^{*}(S^{[n]},\mathbb{Z})$ sending $\epsilon$ to $\delta$. We set
@@
From the classification in Case 1 and Case 2 in Lemma˜7.2, we observe that $\epsilon^{\perp}\cap H^{2}(X,\mathbb{Z})^{\sigma}\neq 0$ or respectively $\epsilon^{\perp}\cap H^{2}(X,\mathbb{Z})^{\sigma-}\neq 0$, so that we are able to pick a non-zero primitive element $\alpha\in\delta^{\perp}\cap H^{2}(S^{[n]},\mathbb{Z})^{\iota-}$ when $\iota(\delta)=\delta$ (Case 2), or $\alpha\in\delta^{\perp}\cap H^{2}(S^{[n]},\mathbb{Z})^{\iota}$ when $\iota(\delta)=-\delta$ (Case 1). Extend $\alpha$ to a basis of $H^{2}(S,\mathbb{Z})$. Consider the basis element
@@
from Theorem˜3.1 and the element
@@
In Case 3 of Lemma˜7.2 (this can only happen when $n\geq 4$ is an even integer): then $A_{L^{\sigma}}\cong\mathbb{Z}/2\mathbb{Z}$ and $A_{L^{\sigma-}}\cong\mathbb{Z}/{(n-1)\mathbb{Z}}$, hence $\tau(\sigma)=-1$. We will use notations as in Section˜4.4. In particular,
@@
Let us choose one $U$-summand in $Q(X,\mathbb{Z})\cong E_{8}(-1)^{\oplus 2}\oplus U^{\oplus 4}$ and pick in it two orthogonal primitive elements $u,v$ such that $u^{2}=2n-2$, $v^{2}=2-2n$ and $u+v$ is divisible by 2n22n-2. According to [67, Theorem 6], there exists an isometry $\phi:Q(X,\mathbb{Z})\to Q(X,\mathbb{Z})$ that sends $u$ to $\gamma$. We put $\delta^{\prime}=\phi(v)$ and define $w:=\frac{\delta^{\prime}+\gamma}{2n-2}\in Q(X,\mathbb{Z})$. In particular, $\delta^{\prime 2}=2-2n$, $\delta^{\prime}\in L^{\prime}$ and the divisibility of $\delta^{\prime}$ in $L^{\prime}$ is 2n22n-2.
@@
For Case 4 of Lemma˜7.2, the proof is the same as in Case 3. The only difference is that now $A_{L^{\sigma}}\cong\mathbb{Z}/(n-1)\mathbb{Z}$ and $A_{L^{\sigma-}}\cong\mathbb{Z}/{2\mathbb{Z}}$, so we have $\tau(\sigma)=1$. Again by Theorem˜4.2, we have the following (instead of (7.8)):
@@ 7.3. Anti-holomorphic involutions
Proof of ˜1.5.
Since maximality is a topological property, we can use hyper-Kähler rotation to change the complex structure. More precisely, let $X$ be a hyper-Kähler manifold of K3[n]-type ($n\geq 2$), and let $\sigma$ be a real structure with respect to the original complex structure $I$. By ˜4.1, there is another complex structure $K$ on $X$ such that $\sigma$ is holomorphic and anti-symplectic. Applying Theorem˜1.4 to the anti-symplectic holomorphic involution $\sigma$ on $(X,K)$, we deduce that $\sigma$ is not maximal. ∎
@@ 8.1. Generalization beyond biregular involutions
By inspecting the proof of Theorem˜1.4, one sees that what is actually proved is the following more general result.
@@
The only place in the proof of Theorem˜1.4 where we used that $\sigma$ is a geometric involution is at the beginning of the proof of Lemma˜7.2, where we argued that the signature of $H^{2}(X,\mathbb{Z})^{\sigma}$ is $(\geq 1,-)$ since $\sigma$ preserves a Kähler class. But the claim for the signature holds in general for any Hodge monodromy operator: since $\sigma\in\operatorname{Mon}^{2}(X)$ is of trivial spinor norm, it preserves $\mathcal{C}^{\circ}_{X}$, the connected component of the positive cone containing Kähler classes. Since $\mathcal{C}^{\circ}_{X}$ is convex, for any Kähler class $\omega$, the class $\frac{\omega+\sigma^{*}(\omega)}{2}$ is $\sigma$-invariant with positive square. Hence the signature of $H^{2}(X,\mathbb{Z})^{\sigma}$ is $(\geq 1,-)$. The rest of the proof of Lemma˜7.2 goes through verbatim. ∎
In addition to Theorem˜1.4 as a special case, Theorem˜8.1 can be applied more generally to birational automorphisms, since a birational automorphism on a compact hyper-Kähler manifold induces a Hodge monodromy operator on its cohomology. Consequently, Theorem˜8.1 implies the following result:
@@ 8.2. Hyper-Kähler varieties of other deformation types
In general, do there exist maximal brane involutions on compact hyper-Kähler manifolds of dimension >2>2? How about the other known deformation types e.g. Kumn-type, OG10-type and OG6-type?
@@
Remark 8.6 (Other varieties with $c_{1}=0$).
Recall that the Beauville–Bogomolov decomposition theorem [3] says that any compact Kähler manifold with vanishing first Chern class admits a finite étale cover that is isomorphic to a product of a complex torus with strict Calabi–Yau manifolds and hyper-Kähler manifolds. Maximal real abelian varieties exist in any dimension [10] [11]. It is expected that maximal real smooth Calabi–Yau hypersurfaces can be constructed by patch-working techniques in any dimension. For the moment complete proofs are available only up to dimension 3, see [14] and the references therein.
8.3. (BBB)-brane involutions on compact hyper-Kähler manifolds
Theorem˜1.4 and ˜1.5 can be reinterpreted as the claim that the Smith inequality Theorem˜1.1 is not optimal for anti-symplectic or anti-holomorphic involutions on compact hyper-Kähler manifolds of K3[n]-type for $n\geq 2$.
Of course, Theorem˜1.6 can be reformulated similarly, but we want to argue that the upper bound given by the Smith inequality for holomorphic symplectic involutions on compact hyper-Kähler manifolds is presumably very far from optimal. Already for K3 surfaces, any symplectic involution has 8 isolated fixed points, whose total $\operatorname{\mathbb{F}_{2}}2$-Betti number is 8, which is much smaller than the one imposed by the Smith inequality (24 in this case).
In fact, we have more precise information on the fixed loci of symplectic involutions on compact hyper-Kähler manifolds of known deformation types. Let $X$ be a compact hyper-Kähler manifold equipped with a holomorphic symplectic involution $\sigma$.
  • If $X$ is of K3[n]-type, by [35, Theorem 1.0.3], the fixed loci of $\sigma$ is the disjoint union of hyper-Kähler manifolds deformation equivalent to Hilbert schemes of points on K3 surfaces [^8] The precise number of connected components in each dimension is given in [35, Theorem 1.0.3].. A computation in the first few values of $n$ shows that the total Betti number of the fixed locus tends to be much smaller than the total Betti number of $X$. For example, when $n=2$, the fixed locus consists of a K3 surface and 28 points, whose total Betti number is 52; when $n=3$, the fixed locus consists of 8 K3 surfaces and 64 points, whose total Betti number is 256; when $n=4$, the fixed locus consists of 126 points, 28 K3 surfaces and a hyper-Kähler manifold of K3[2]-type, whose total Betti number is 1122. The manifold $X$ has total Betti number 324 for $n=2$, 3200 for $n=3$ and 25650 for $n=4$.
  • If $X$ is of $\operatorname{Kum}_{n}$-type, some partial results are obtained in [35, Theorem 1.0.4].
  • If $X$ is of OG6-type, by [25, Theorem 1.1], $\sigma$ acts trivially on the second cohomology. Such involutions are classified in [51, Theorem 5.2], and their fixed loci are always disjoint unions of K3 surfaces and isolated points.
  • If $X$ is of OG10-type, by [22, Theorem 1.1], there is no non-trivial finite-order symplectic automorphism.
Question 8.7.
@@
We provide a proof of Theorem˜1.6 covering also OG6-type hyper-Kähler manifolds. The key observation is the following elementary fact in topology:
Lemma 8.8.
Let $X$ be a compact manifold equipped with an involution $\tau$ acting non-trivially on $H^{*}(X,\mathbb{Q})$. Assume $H^{\operatorname{odd}}(X,\mathbb{Q})=0$ and $H^{\operatorname{odd}}(X^{\tau},\operatorname{\mathbb{F}_{2}}2)=0$. Then $\tau$ is not maximal.
Proof.
Let $b_{*}(-,\operatorname{\mathbb{F}_{2}}2)$ denote the total $\operatorname{\mathbb{F}_{2}}2$-Betti number.
$b_{*}(X^{\tau},\operatorname{\mathbb{F}_{2}}2)$
$=$ $\chi(X^{\tau})$
$=$ $2\chi(X/\tau)-\chi(X)$
$=$ $\sumop\displaylimits_{i}\left(2\dim H^{i}(X,\mathbb{Q})^{\tau}-\dim H^{i}(X,\mathbb{Q})\right)$
$<$ $\sumop\displaylimits_{i}\dim H^{i}(X,\mathbb{Q})$
$\leq$ $b_{*}(X,\operatorname{\mathbb{F}_{2}}2)$
where the first equality is by the hypothesis that $H^{\operatorname{odd}}(X^{\tau},\operatorname{\mathbb{F}_{2}}2)=0$, the third equality is by the vanishing of $H^{\operatorname{odd}}(X,\mathbb{Q})$, the strict inequality follows from the assumption that $\tau$ acts non-trivially on $H^{*}(X,\mathbb{Q})$, the last inequality is from the universal coefficient theorem. ∎
Corollary 8.9.
A non-trivial holomorphic symplectic involution on a compact hyper-Kähler manifold of K3[n]-type or OG6-type is not maximal.
Proof.
It suffices to check the conditions of Lemma˜8.8. By [2, Proposition 10] for K3[n]-type, [51, Theorem 5.2 and Remark 6.9] for OG6-type, there is no nontrivial automorphisms acting trivially on $H^{*}(X,\mathbb{Q})$. By [23] for K3[n]-type, [50] for OG6-type, $H^{\operatorname{odd}}(X,\mathbb{Q})=0$. By [35, Theorem 1.0.3] for K3[n]-type, [25] and [51] for OG6-type, the fixed locus $X^{\tau}$ is a disjoint union of hyper-Kähler manifolds of K3[m]-type (for various $m$), hence $H^{\operatorname{odd}}(X^{\tau},\operatorname{\mathbb{F}_{2}}2)=0$. ∎